Iron carbonyl can be made by the direct reaction of iron metal and carbon monoxide.
Fe(s)+5 CO(g)--->Fe(CO)5(l)
What is the theoretical yield of Fe(CO)5 if 3.01 g of iron is treated with CO gas having a pressure of 732 mm Hg in a 5.36-L flask at 23 °C?

Theoretical yield = ?? g
1
Expert's answer
2020-05-03T15:14:27-0400

m (Fe) = 3.01 g

n(Fe) = m / M = 3.01 g / 55.845 g/mol = 0.054 mol

P(CO) = 732 mm Hg / 760 mm Hg = 0.963 atm

T(CO) = 23°C = 23 + 273.15 = 296.15 K

V(CO) = 5.36 L

P x V = n x R x T

n = P x V / R x T = 0.963 atm x 5.36 L / 0.082057 L•atm/(K•mol) x 296.15 K = 0.212 mol

Fe(s) + 5CO(g) ---> Fe(CO)5(l)

Fe is excess reactant and CO is limiting rectant

n(Fe(CO)5) = n(CO) / 5 = 0.212 / 5 = 0.0424 mol

m(Fe(CO)5) = n(Fe(CO)5) x M(Fe(CO)5) = 0.0424 mol x 196 g/mol = 8.31 g